Two insulating small spheres are rubbed against each other and placed 1 cm apart, If they attract each other with a force of 0.1 N, how many electrons were transferred from one sphere to the other during rubbing?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. F = 0.1 N r = 1 cm = 10^–2 (As they rubbed with each other. So the charge on each sphere are equal) So, F = kq base 1q base 2/r^2 ⇒ 0.1 = kq^2/(10^-2)^2 ⇒ q^2 = 1/9 * 10^-14 ⇒ q = 1/3 * 10^-7 1.6 * 10^-19 c Carries by 1 electron 1 c carried by 1/1.6 * 10^-19 0.33 * 10^-7 c carries by 1/1.6 * 10^-19 * 0.33 * 10^-7 = 0.208 * 10^12 = 2.08 * 10^11